Grade 11Mechanics

Water from a pipe of cross sectional area 10^-2 m^2 gushes out of the tube horizontal ly with a speed of 15m/sec .The water hits a vertical wall nearby and rebounds with the same speed What is the force exerted by the impact of water.

Paridhi Chawhan
6 Years agoGrade 11

1 Answer

Vikas TU
6 Years ago
Cross sectional area = 10^-2
Spped = 15m/sec 
Vol of water coming out from pipe 
V = Av = 15 * 10^-2
Density of water = 10^3 kg/m^3
Mass of water flowing out through pipe = row*v = 150 kg /sec
Force extended by water on wall =Rate of change of momentum = delP /del T 
= mv / t 
=> 150*15 = 2250N